OpenStudy (anonymous):

how do I solve the 7th term of the sequence of 1,4,16,64?

OpenStudy (anonymous):

Notice how each term is being multiplied by 4 each time? This is called a geometric sequence which has the formula a(n)=a(1)*r^(n-1) where a(1) is the first term of the sequence, r is the common ratio, n is the number of the term to find.

OpenStudy (anonymous):

So let's say we want the 7th term. a(7)=a(1)*r^(7-1)

OpenStudy (anonymous):

r is our common ratio, or what we multiply by every time, which is 4. a(7) is what we are looking for a(1) is our first term and is 1 a(7)=1*4^6 a(7)=4096 Hope this helped!

OpenStudy (anonymous):

Is there a better answer?

OpenStudy (anonymous):

How can I solve the 7 term of sequence of 1,4,16,64 in a understandable step by step way?

OpenStudy (anonymous):

The easiest way for me to explain this is look at the sequence. You want the 7th term of this sequence. Notice any patterns between each number?